Test: Test 2 Ch 5-6 Time Remaining: 025309 Submit Test This 1 pt This Test: 30 pts does, fnd the mean and standard deviation 001650104902671 0 3400 OA. μ. (Round to one decimal place as needed) O B. The table is not a probability disaribution Fnd the standard deviation of the random variable x Select the comect choice below and, if necessary, ill in the answer box to complete your choice ○A.。“ (Rond to one decamal pace as needed ) O B. The table is not a probabiity distribution 丱@ DOLL
media%2F996%2F996a4c38-0d9a-4fcb-939e-44
Answer #1

Solution:

Mean = \mu = \sum x.P(x)

= 0*0.0165 + 1*0.1049 + 2*0.2671 + 3*3400 + 4*2164 + 5*0.0551

= 2.8002

  \approx 2.8

Standard deviation = \sigma = \sqrt{\sum x^{2}p(x)-\mu ^{2}}

Here,

\sum x^{2}p(x) =  02*0.0165 + 12*0.1049 + 22*0.2671 + 32*3400 + 42*2164 + 52*0.0551

= 9.0732

\sigma = \sqrt{\sum x^{2}p(x)-\mu ^{2}} = \sqrt{9.0732-2.8^{2}}

Standard deviation = \sigma =1.1